85 inches to feet to yard to meter
Found 2 solutions by rothauserc, Theo:
Answer by rothauserc(4718)

85/12 = 7.08 feet
7.08 / 3 = 2.36 yards
2.36 * 0.9144 = 2.16 meters

Answer by Theo(13342)

i'll round everything to 6 places.
this should get you an answer accurate to 4 decimal places.

use a conversion table to find feet per inch, yards per feet, meters per yard.

the conversion table will tell you:

1 inch = 1/12 feet.
1 foot = 1/3 yards.
1 yard = .9144 meters.

if 1 inch = 1/12 feet, then 85 inches = 85 * 1/12 = 85/12 feet.

if 1 foot = 1/3 yards, then 85/12 feet = 85/12 * 1/3 = 85/36 yards.

if 1 yard = .9144 meters, then 85/36 yards = 85/36 * .9144 = 77.724/36 = 2.159 meters.

you can check your answer by working backwards from meters to yards to feet to inches.

check your conversion tables again:

my conversion tables showed me:

1 meter = 1.09361 yards
1 yard = 3 feet
1 foot = 12 inches

if 1 meter = 1.09361 yards, then 2.159 meters = 2.159 * 1.09361 = 2.361104 yards.

if 1 yard = 3 feet, then 2.361104 yards = 2.361104 * 3 = 7.083312 feet.

if 1 foot = 12 inches, then 7.083312 feet = 7.083312 * 12 = 84.999744 inches.

the answer isn't right on because of rounding.

the most likely culprit is the conversion from meters to yards and yards to meters.

the conversion table lists 1 yard = .9144 meters.
the conversion table lists 1 meter = 1.09361 yards.

if 1 meter = 1.09361 yards, then divide both sides of this equation by 1.09361 to get:
1 meter / 1.09361 = 1 yard.
simplify to get .914403 meters = 1 yard.

the conversion table said that 1 yard = .9144 meters.
there was some rounding being done.
this threw off the confirmation.
we got close to 85 inches but not right on.
if there was no rounding, we would have been right on.